OpenStudy (anonymous):

sqrt{125^x} = 5/25^x

OpenStudy (anonymous):

hm?

OpenStudy (anonymous):

thats alright. want me to put my other questions here?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{125^x} = \frac{5}{25^x}\] \[\Large \sqrt{\left(5^3\right)^x} = \frac{5^1}{\left(5^3\right)^x}\] \[\Large \sqrt{5^{3x}} = \frac{5^1}{5^{2x}}\] \[\Large \left(5^{3x}\right)^{\frac{1}{2}} = \frac{5^1}{5^{2x}}\] \[\Large 5^{\frac{3x}{2}} = \frac{5^1}{5^{2x}}\] \[\Large 5^{\frac{3x}{2}} = 5^{1-2x}\] \[\Large \frac{3x}{2} = 1-2x\] \[\Large \frac{3x}{2}+2x = 1\] \[\Large \frac{7x}{2} = 1\] \[\Large 7x = 2\] \[\Large x = \frac{2}{7}\]